If 4.24 g of mixture of NaCl and Na2CO3 is reacted completely with 25 ml of 0.4 m H2SO4 solution then the percentage of NaCl in the mixture is
a. 32 
b. 75
c. 62
d. 45

Grade:12

1 Answers

Vikas TU
4 years ago
a+b = 4.24 g 
In 25 ml Meq of Na2Co3 = Meq of Hcl 
= 50 * 1/10
= 5 
Meq in 250 ml = 50 
a/ (106/2)  * 1000 = 50 
a = 2.65 
b = 4.24 - 2.65 = 1.59 
% Nacl = 32% 
Hope this helps